OpenStudy (rational):

base conversion \[\large (0.\overline{1210})_3 = (?)_{10}\]

OpenStudy (inkyvoyd):

oops looks like my logic is right by my arithmetic was wrong

OpenStudy (inkyvoyd):

from there you can apply the same logic to base 2 to get .1210 repeating= 1210_3/2222_3 = 48_10/80_10=3/5=.6

OpenStudy (rational):

wow is that a shortcut XD i thought u would expand it and use geometric series formula

OpenStudy (inkyvoyd):

@rational I am so glad my algebra teacher taught me this trick a long time ago lol

OpenStudy (inkyvoyd):

it is basically geometric series but presented in a way that is more simple than it is normally taught

OpenStudy (rational):

That looks neat! so for a repeating pattern of length \(r\), do we have \[\large (0.\overline{abcd\ldots })_b = \dfrac{(abcd\ldots)_b}{(b^r-1)_{10}}\]

OpenStudy (inkyvoyd):

yeah, I learned it as a rule a long time ago, and we used algebra to prove it, for the decimal case at least... I figured that the logic was behind radix representation of numbers, which is why it would just as readily apply to any base as well

OpenStudy (inkyvoyd):

the base 10 should actually be b I believe... so you would see (call c=b-1) ccccc... for the length of abcd...

OpenStudy (rational):

\[(ccc\ldots )_{b} = (c(111\ldots))_b \\~\\= (c(b^{r-1} + b^{r-2}+\cdots + b^1+1))_{10} \\~\\= \left(c\dfrac{b^r-1}{b-1}\right)_{10} \\~\\ = (b^r-1)_{10}\]

OpenStudy (inkyvoyd):

hahaha oh